OpenStudy (rissyroo118):

x + 3y = 5 -x + 6y = 4 Solve the system of equations. A) x = 1, y = 2 B) x = 2, y = 1 C) x = 1, y = 1 D) x = 0, y = 2 @Leader

OpenStudy (laweejay):

B

OpenStudy (rissyroo118):

A) (-1, -3) B) (1, -3) C) (1, 3) D) (3, 1)

OpenStudy (laweejay):

C

OpenStudy (rissyroo118):

A) y = x + 2 and y = -x - 4 B) y = 3x + 7 and y = x + 12 C) y = 2x - 3 and y = -2x + 1 D) y = -x + 4 and y = -2x + 5

OpenStudy (leader):

9y=9 y=1 x=2 So it is B not C

OpenStudy (leader):

And 1 question per post please.

OpenStudy (rissyroo118):

okay :) my 2nd question was not c it was b correct ?

OpenStudy (laweejay):

*the C was for the 2nd question...

OpenStudy (leader):

Your fist question is B